# Model Context Protocol (MCP): Landscape, Security Threats, and Future Research Directions

### v2

@@ 7.1. Tool Integration in LLM Applications
As tool-augmented LLMs continue to evolve, the lack of a standardized, secure, and extensible context protocol has become a key bottleneck. MCP, with its potential to unify tool interaction across diverse systems, is poised to become the foundational layer for next-generation LLM applications, making it critical to examine its landscape, limitations, and risks.
@@ 7.2. Security Risks in LLM-Tool Interactions
These security threats may be mitigated through the structured design of MCP, but they can also persist or even evolve under this new integration paradigm. As MCP simplifies tool orchestration in LLM applications, it simultaneously introduces new potential attack surfaces, warranting deeper investigation into its security implications.
@@ 8. Conclusion
This paper presents the first comprehensive analysis of the MCP ecosystem landscape. We examine its architecture, core components, operational workflows, and server lifecycle stages. Furthermore, we explore the adoption, diversity, and use cases, while identifying potential security threats throughout the creation, operation, and update phases. We also highlight the implications and risks associated with MCP adoption and propose actionable recommendations for stakeholders to enhance security and governance. Additionally, we outline future research directions to tackle emerging risks and improve MCP’s resilience. As MCP continues to gain traction with industry leaders such as OpenAI and Cloudflare, addressing these challenges is key to its long-term success and to enabling secure, efficient interaction with diverse external tools and services.

### v3

@@ 7.1. Tool Integration in LLM Applications
As the demand for LLMs to interact with external tools continues to grow, the research community has increasingly focused on improving the quality, reliability, and generality of such interactions. Existing work demonstrates clear progress, from automatic tool discovery and documentation parsing to multi-step reasoning and evaluation, but these efforts remain fragmented and platform-specific. The lack of a unifying, secure, and extensible protocol leads to duplicated engineering, inconsistent integration practices, and heightened security risks. This fragmentation highlights the value and necessity of the MCP. MCP is designed to provide a standardized infrastructure for LLM–tool communication, defining unified formats for context exchange, capability invocation, and access control. Given its emerging role at the core of tool-augmented LLM ecosystems, it becomes crucial to comprehensively map MCP’s current landscape, analyze its security implications, and discuss its future evolution. This investigation provides both the conceptual grounding and practical guidance needed to foster a safer, more standardized foundation for LLM–tool integration.
@@ 7.2. Security Risks in LLM-Tool Interactions
The above studies collectively reveal the multifaceted security risks inherent in LLM–tool interaction, spanning prompt manipulation, unsafe execution, and untrusted endpoints. These findings motivate a deeper examination of how such risks may evolve under the emerging MCP. It remains unclear whether the structured design of MCP will merely encapsulate these existing vulnerabilities, exacerbate them through new cross-context channels, or provide the means to mitigate them through standardized control and isolation. Understanding this shift is central and constitutes a primary motivation of this work, as it is essential for evaluating the true security implications of MCP within future tool-augmented LLM ecosystems.
7.3. Security of the MCP
As the MCP becomes a foundational interface for tool-augmented AI ecosystems, its security and reliability have attracted increasing attention. Its open and flexible design enables broad interoperability but also introduces novel and systemic security risks. Recent studies have primarily focused on characterizing these risks and identifying vulnerabilities. Hasan et al. (Hasan et al., 2025) conduct the large-scale empirical measurement of 1,899 open-source MCP servers, revealing eight novel vulnerability categories arising from immature maintenance and non-deterministic control flow. Zhao et al. (Zhao et al., 2025b) provide a systematic taxonomy of malicious MCP server behaviors and demonstrate the feasibility of practical attacks through proof-of-concept exploits. Wang et al. (Wang et al., 2025b) describe the Preference Manipulation Attack (MPMA), in which adversarial servers bias LLM decisions for economic gain within open marketplaces. Similarly, Shuli Zhao et al. (Zhao et al., 2025a) identify Parasitic Toolchain Attacks that exploit weak context isolation, resulting in stealthy data exfiltration across interconnected tools. To standardize empirical testing, Yang et al. (Yang et al., 2025) introduce MCPSecBench, which defines a benchmark of seventeen attack types across multiple host and client configurations, enabling reproducible MCP security assessments. Radosevich and Halloran (Radosevich and Halloran, 2025) further highlight that existing MCP workflows allow severe code execution and privilege escalation exploits, revealing insufficient safeguards in current implementations.
Complementary work explores proactive protection and hardening mechanisms for the MCP ecosystem. Bhatt et al. (Bhatt et al., 2025) propose the Enhanced Tool Definition Interface (ETDI), integrating OAuth-based identity verification and policy-based access control to mitigate squatting and rug-pull attacks. Xing et al. (Xing et al., 2025) develop MCP-Guard, a layered defense architecture combining static signature scanning, deep neural threat detection, and LLM-based decision arbitration. Kumar et al. (Kumar et al., 2025) present MCP Guardian, a defense-in-depth layer that strengthens authentication, rate limiting, logging, and firewall integration for MCP communications. At the enterprise level, Brett (Brett, 2025) introduces MCP Gateways to simplify secure self-hosted deployments through threat model mapping, authentication, and intrusion detection, supporting enterprise-grade isolation.
Most prior studies focus on isolated phases, either vulnerability scanning or runtime protection, without establishing a lifecycle, wide security model. They also tend to emphasize specific attack vectors or detection accuracy. In contrast, this work provides the first end-to-end security and privacy analysis of the MCP ecosystem covering all lifecycle stages. This lifecycle-centric analysis bridges micro-level vulnerabilities and macro-level protocol design, enabling the derivation of practical mitigation strategies and secure-by-design guidelines for the future evolution of the MCP.
@@ 8. Conclusion
This paper presents the first comprehensive analysis of the MCP ecosystem landscape. We examine its architecture, core components, operational workflows, and server lifecycle stages. Furthermore, we explore the adoption, diversity, and use cases, while identifying potential security threats throughout the creation, deployment, operation, and maintenance phases. We also highlight the implications and risks associated with MCP adoption and propose actionable recommendations for stakeholders to enhance security and governance. Additionally, we outline future research directions to tackle emerging risks and improve MCP’s resilience. As MCP continues to gain traction with industry leaders such as OpenAI and Cloudflare, addressing these challenges is key to its long-term success and to enabling secure, efficient interaction with diverse external services.